# Deep Emotion Recognition in Textual Conversations: A Survey

### v3

@@ 4.7.1 Conversational Memory Network [37]
The recurrent networks described before are limited in terms of long-range summarization. The Conversational Memory Network (CMN) accounts for this factor by the use of memory networks, which can better capture long-term dependencies and have attention models that can summarize specific details.
The work is based on the notion that emotional dynamics involves self and inter-speaker emotional influence. Separate histories for each speaker, consisting in the speaker’s previous utterances, are modeled into memory cells of the Conversational Memory Network using GRUs. Memory networks provide a memory component that can be read from and written to and also perform inference. The CMN then employs an attention mechanism over historical utterances from each speaker to filter out relevant content for the current utterance. This mechanism is repeated for multiple hops of the network to subsequently classify the utterance. CMN achieves an F1-score of 56.13% in the IEMOCAP dataset.
@@ 4.8 Graph Neural Networks
An alternative to gated neural networks that is also useful to capture long term dependencies is using graph neural networks. Graph neural networks are based on graphs. The latter model a set of objects, nodes, and their relationships, edges. While standard neural networks such as CNNs and RNNs cannot handle the graph input properly since they need a specific order for the nodes, the output of GNNs is invariant for the input order of nodes. Furthermore, while in standard neural networks, dependencies between nodes are a node feature, in GNNs there are edges to represent these dependecies. GNNs can propagate information by the graph structure instead of using it as part of features [150].
Two common types of GNNs are recurrent GNNs (RecGNNs) and convolutional GNNs (ConvGNNs). RecGNNs learn node representations with recurrent neural architectures, assuming that a node in a graph constantly exchanges information with its neighbors until a stable equilibrium is reached. ConvGNNs generalize the convolution operation from grid structure to graph structure, generating a node representation by aggregating its features with its neigbours’ features. For classification tasks, such as Emotion Recognition, an end-to-end framework can be constructed for example by stacking graph convolutional layers followed by a softmax layer [131].
@@ 4.8.1 DialogueGCN [31]
Dialogue Graph Convolutional Network (GCN) takes into account intra and inter-speaker dependency and background information, having a sequential and a speaker-level encoder. The sequential context encoder encodes the utterances using a bidirectional GRU and is speaker agnostic. The speaker-level context encoder creates a directed edge-labeled graph in which each utterance is a node in the graph. The informations from neighbor nodes are then aggregated and passed to a neural network in each node, resulting in an updated representation of each node that takes into account its neighbors. In DialogueGCN, node features are initialized with sequentially encoded feature vectors from the sequential context encoders and edge weights are set using a similarity based attention module. The utterance level emotion classification is turned into a problem of node classification in the graph. DialogueGCN achieves an F1-score of 64.18% on the IEMOCAP dataset.
@@ 4.9 Transformer
Despite the success of recurrent and graph neural networks in a wide variety of tasks including Emotion Recognition, a more powerful network model exists, constituting the current state of-the-art: the Transformer [120]. The Transformer is also better at capturing long-term dependencies than gated RNNs, due to its shorter path of information flow, which is useful when modelling several utterances in Emotion Recognition in Conversations. It does not resort to recurrence nor convolutions. It is also more parallelizable and needs less training time. It is composed of an Encoder and Decoder, that can be visualized in Figure 7.
@@
The self-attention mechanism takes as input three matrices, the Query, Key and Value matrices. These are computed from the dot product between the input sequence matrix and weight matrices learned in the training phase. It then computes a vector of attention scores, as described in the following equation:
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{k}})V,$ (15)
in which $Q\in\mathbb{R}^{N\times m}$, $K\in\mathbb{R}^{N\times m}$ and $V\in\mathbb{R}^{N\times m}$ are the Query, Key and Value matrices, respectively, and $m$ is the model output dimension. The dot product of the Query matrix with the Key matrix represents scores that are then multiplied by the Value matrix to keep intact the values of relevant words and drown-out the irrelevant ones. The $\frac{1}{\sqrt{k}}$ helps stabilizing the gradients.
@@
Each Decoder layer is composed of three sublayers. The first is a masked self-attention layer similar to the self-attention layer described mathematically in Equations 16 and 17, but in which each position in the decoder attends to all positions in the decoder up to and including that position, which is attained by masking out values in the softmax. The second is a “encoder-decoder attention” layer, described mathematically in the same equations, where K and V come from the output of the encoder and Q comes from the output of the decoder’s masked self-attention layer, mimicking the typical encoder-decoder attention mechanisms in Seq2Seq models. The last layer is a feed-forward neural network, similar to the Encoder layer.
@@
The most common use of the Transformer is as backbone for a fine-tune pre-trained transformer-based language model, such as BERT [20], described in subsubsection 4.9.4. These language models are trained with large amounts of data to perform many tasks and only need to be fine-tuned for the specific task at hand. In the case of a classification task such as Emotion Recognition one just needs to add a linear layer or a more complex neural network architecture on top of the language model. The training is performed with supervised learning, resorting to backpropagation algorithms.
@@ 4.9.1 Knowledge-Enriched Transformer [149]
The invention of the Transformer [120] led to new state-of-the-art results in several Natural Language Processing tasks. The Knowledge-Enriched Transformer (KET) uses self-attention to model context and response using the Transformer, that a has shorter path of information flow than gated RNNs, overcoming the difficulty in capturing long-term dependencies. Conversations are modeled in the entire Transformer.
First, concepts are retrieved for each word in the conversation using an external knowledge base, a graph of concepts, then the concept representation is computed using a dynamic context-aware graph attention mechanism over the concepts, using measures of relatedness and affectiveness. Self-attention is used to learn utterance representations individually and hierarchical self attention is applied to the context to learn the context representation. After having the context passed into the encoder the response is passed into the decoder. Finally, the encoder and decoder attention mechanism is applied, followed by max pooling and a linear layer. KET achieves an F1-score of 59.56% on the IEMOCAP dataset.
@@ 4.9.3 DialogXL [107]
DialogXL is an adaptation of the pretrained language model XLNet [137], which is based on the Transformer-XL [15], for Emotion Recognition in Conversations. It replaces XLNet’s segment recurrence by a memory utterance recurrence to leverage historical utterances. These utterances’ hidden states are stored in a memory bank to reuse them while identifying a query utterance. The approach also replaces XLNet’s vanilla self-attention proposing a dialog-aware self-attention to grasp useful intra- and inter-speaker dependencies. This is composed of four types of self-attention: global and local self-attention for different sizes of receptive fields, and speaker and listener self-attention for intra- and inter-speaker dependencies. DialogXL achieves an F1-score of 65.94% on the IEMOCAP dataset.
4.9.4 Embeddings from Pre-Trained Transformer Language Models
@@
Static embeddings [83][90][84] are obtained based on the co-occurrence of adjacent words and have a fixed representation for the words, not taking into account its context.
Contrarily, contextual embeddings include information from the context in the word representation, a representation that differs according to the occurrence of the word. The only disadvantage of this kind of representation is that we cannot use pre-trained embeddings as in static embeddings so there is an inference time for generating the embeddings. The models are pre-trained with large amounts of data, in order to enable generating good embedings. A first approach, Embeddings from Language Models (ELMo) [92] is based on a stack of two Bi-LSTMs, leveraging the full context of a word. It provides a context free representation of the word along with context information of the sense of the word and its syntax. The second approach, Bidirectional Encoder Representations from Transformers (BERT) [20], built from a Transformer encoder stack, leads to state-of-the art results in multiple Natural Language Processing tasks. The input given to BERT is the sum of three embeddings, the token embedding corresponding to the sequence of words, the segment embedding containing information on which sentence each token belongs to, and the position embedding, defining the position of each token in the sequence and distance to other tokens taking into account the order of the sequence. An improvement upon BERT, RoBERTa [73] is trained with more data and for a longer period of time. Given that context is key for text Emotion Recognition in Conversations, contextual word embeddings are more adequate for the task than static embeddings, as it can be seen in the higher performance of works that resort to these embeddings, namely resultant from Pre-trained language models, such as BERT-based embeddings. The latter are also more tailored to deal with common sense, informal language and sarcasm, ERC challenges described in Section 2, since they are pre-trained on a large amounts of text.
@@ 4.9.5 COSMIC [30]
COSMIC extracts commonsense features resorting to COMET [5], a commonsense transformer encoder-decoder model trained on the task of generative commonsense knowledge construction that uses GPT [101] as its generative model. The approach uses RoBERTa for context independent feature extraction, passes each uterrance concatenated with a relation phrase through COMET’s encoder and extracts the activations from the final time-step. COSMIC maintains a context state and attention vector which are always shared between the participants of the conversation. Five GRUs are used to model context state, internal state, external state, intent state, and emotion state, being the latter used for the emotion classification. This work achieves an F1-score of 65.28% in the IEMOCAP dataset.
@@ 4.9.6 Psychological [65]
Psychological also leverages COMET as commonsense knowledge base and resorts to RoBERTa as utterance-level encoder. It proposes SKAIG as a conversation-level encoder, a locally connected graph, where the targeted utterance receives information from that past and future context and is also self-connected.
Assuming that the influence of an utterance on contextual utterances is locally effective, the targeted node is connected with contextual nodes in a window of a given size of utterances in the past and future. Knowledge from COMET is introduced to enrich the edges with diferent relations. To propagate information through SKAIG a Graph Transformer [110] is used. Finally a linear unit predicts the emotion distributions. This work achieves an F1-score of 66.96% in the IEMOCAP dataset.
@@ 4.9.7 DAG-ERC [108]
Graph-based methods tend to neglect the distant utterances and sequential information. On the other hand, recurrence-based methods leverage those, but tend to update the query utterance’s state with limited information from the nearest utterances. Combining both graph and recurrence structures explores their advantages and mitigates their drawbacks. This is the idea behind DAG-ERC, that regards each conversation as a directed acyclic graph (DAG), a combination of both structures. DAG-ERC is based on the the DAGNN [113] architecture, with two improvements: a relation-aware feature transformation that gathers information based on speaker identity and a contextual information unit that enhances the information of historical context. The state-of-the art RoBERTa-Large is used as feature extractor. It achieves an F1-score of 68.03% in the IEMOCAP dataset, outperforming the previous approaches.
4.9.8 ERC in a Generative Framework
With the advent of Generative Large Language Models (LLMs), such as the GPT series [100][101][7] and Llama [114], classification tasks started to be reformulated within generative frameworks. The idea is to prompt the LLM with the utterance to be classified and other relevant additional input, and request an emotion label for the utterance. Open source LLMs can also be fine-tuned and even pre-trained.
@@ 4.9.9 InstructERC [62]
In InstructERC the prompt consists in the utterance to be classified, the ERC instruction, the conversational context, the set of possible classification labels and similar utterances together with its emotion labels. The LLM is pre-trained with a speaker identification task and fine-tuned both with the main ERC task and an emotion influence prediction task to improve overall performance. Experiments were made with several LLMs, and best results were obtained with LLama2. InstructERC achieves 71.39% in the IEMOCAP dataset, outperforming the previous approaches.
@@ 5 Advisable ERC Practices
This section provides useful methods towards better ERC frameworks, from different ways of dealing with subjectivity in emotion annotation and modelling, to several ways of dealing with the typicaly unbalanced ERC datasets.
@@ 5.1 Learning Subjectivity
All human annotation tasks involve some degree of uncertainty, that can stem amongst other factors from the subjectivity inherent to the annotator, being Emotion Recognition amongst the most subjective tasks [119]. Given that subjectivity is the main source of uncertainty in Emotion Recognition annotation, we address uncertainty as subjectivity in this survey [104]. A good annotation practice is to resort to several annotators to avoid biasing the results towards a single annotator’s subjectivity. This will naturally yield different labels from each annotator, the so called annotation disagreement. The more appropriately one deals with the disagreement in the annotated labels the more reliable will the data and classifiers be. Several techniques for dealing with disagreement are described in this section. They are divided into techniques that address subjectivity by adapting the labels, towards an estimate of a gold label, and techniques that embrace subjectivity by adapting the classifier architectures to deal with the subjectivity of the labels, leveraging such property.
@@ 5.1.1 Addressing Subjectivity
For categorical labels, considering that a gold or true label exists for each sample, the simple way to obtain it would be through majority voting, choosing the label with the most annotations. However, this does not account for the different annotator skill levels neither sample inherent difficulty.
@@
For interval labels, weighted fusion of data can be performed with approaches such as the Evaluator Weighted Estimator (EWE) [34], that weights each rater’s annotations based on its rater-specific inter-reliability scores, or the Weighted Trustability Evaluator (WTE) [35], which is instead based on each rater’s performance consistency.
@@ 5.1.2 Embracing Subjectivity [104]
The annotator (dis)agreement can be used as privileged information for master-class learning, being viewed as samples’ additional information [25] to facilitate learning in the training phase, that might or not be required for making predictions during testing to, for example, weigh positively the loss of the corresponding samples. In this setting it is useful to learn the meaning of high rater disagreement means for the particular dataset in use.
Instead of considering the existence of a hard label, as in subsection 5.1.1, one can calculate the distribution of label annotations to define a soft label distribution per sample, which encodes the subjectivity for each sample and allows for the classifier to learn label correlations. Another approach is to model the annotations of each particular rater, through an ensemble of models. It allows to give more importance to expert raters rather than novices or spammers. The addition of an "unsure" label would facilitate annotator’s workload and also make the model learn the ambiguity ground-truth, although not having to predict "unsure" labels during testing. Furthermore, methods to understand whether a sample is mislabelled by certain raters or inherently ambiguous could inform the model and annotators during active learning processes.

### v4

@@ 4.7.1 Conversational Memory Network [31]
The recurrent networks described before are limited in terms of long-range summarization since they rely on a sequential processing approach with implicit memory. The Conversational Memory Network (CMN) accounts for this factor through the use of memory networks, which can better capture long-term dependencies and have attention models that can summarize specific details using explicit memory structures.
The work is based on the notion that emotional dynamics involve self and inter-speaker emotional influence. Separate histories for each speaker, consisting of the speaker’s previous utterances, are modeled into memory cells of the Conversational Memory Network using GRUs. Memory networks provide a memory component that can be read from and written to and also perform inference. The CMN then employs an attention mechanism over historical utterances from each speaker to filter out relevant content for the current utterance. This mechanism is repeated for multiple hops of the network to subsequently classify the utterance. CMN achieves an F1-score of 56.13% on the IEMOCAP dataset.
@@ 4.8 Graph Neural Networks
An alternative to gated neural networks that is also useful to capture long term dependencies is using graph neural networks. Graph neural networks are based on graphs. The latter model a set of objects, the nodes, and their relationships, the edges. While standard neural networks such as CNNs and RNNs cannot handle the graph input properly since they need a specific order for the nodes, the output of GNNs is invariant for the input order of nodes. Furthermore, while in standard neural networks, dependencies between nodes are a node feature, in GNNs there are edges to represent these dependencies. GNNs can propagate information by the graph structure instead of using it as part of features [145].
Two common types of GNNs are recurrent GNNs (RecGNNs) and convolutional GNNs (ConvGNNs). RecGNNs learn node representations with recurrent neural architectures, assuming that a node in a graph constantly exchanges information with its neighbors until a stable equilibrium is reached. ConvGNNs generalize the convolution operation from grid structure to graph structure, generating a node representation by aggregating its features with its neighbours’ features. For classification tasks, such as Emotion Recognition, an end-to-end framework can be constructed for example by stacking graph convolutional layers followed by a softmax layer [122].
@@ 4.8.1 DialogueGCN [26]
Dialogue Graph Convolutional Network (GCN) takes into account intra and inter-speaker dependency and background information, having a sequential and a speaker-level encoder. The sequential context encoder encodes the utterances using a bidirectional GRU and is speaker-agnostic. The speaker-level context encoder creates a directed edge-labelled graph in which each utterance, enriched with context, is a node in the graph. The information from neighbor nodes is then aggregated and passed to a neural network in each node, resulting in an updated representation of each node that takes into account its neighbors. In DialogueGCN, node features are initialized with sequentially encoded feature vectors from the sequential context encoders, and edge weights are set using a similarity-based attention module. The utterance level emotion classification is turned into a problem of node classification in the graph. DialogueGCN achieves an F1-score of 64.18% on the IEMOCAP dataset.
@@ 4.9 Transformer
Despite the success of recurrent and graph neural networks in a wide variety of tasks including Emotion Recognition, a more powerful network model exists, constituting the current state-of-the-art: the Transformer [112]. The Transformer is also better at capturing long-term dependencies than gated RNNs, due to its shorter path of information flow, which is useful when modelling several utterances in Emotion Recognition in Conversations. It does not resort to recurrence or convolutions. It is also more parallelizable and needs less training time. It is composed of an Encoder and Decoder, which can be visualized in Figure 7.
@@
The self-attention mechanism takes as input three matrices, the Query, Key, and Value matrices. These are computed from the dot product between the input sequence matrix and weight matrices learned in the training phase. It then computes a vector of attention scores, as described in the following equation:
$Attention(Q,K,V)=softmax\left(\frac{QK^{T}}{\sqrt{k}}\right)V,$ (15)
in which $Q\in\mathbb{R}^{N\times m}$, $K\in\mathbb{R}^{N\times m}$ and $V\in\mathbb{R}^{N\times m}$ are the Query, Key, and Value matrices, respectively, and $m$ is the model output dimension. The dot product of the Query matrix with the Key matrix represents scores that are then multiplied by the Value matrix to keep intact the values of relevant words and drown-out the irrelevant ones. The $\frac{1}{\sqrt{k}}$ helps stabilizing the gradients.
@@
Each Decoder layer is composed of three sublayers. The first is a masked self-attention layer similar to the self-attention layer described mathematically in Equations 16 and 17, but in which each position in the decoder attends to all positions in the decoder up to and including that position, which is attained by masking out values in the softmax. The second is an “encoder-decoder attention” layer, described mathematically in the same equations, where K and V come from the output of the encoder and Q comes from the output of the decoder’s masked self-attention layer, mimicking the typical encoder-decoder attention mechanisms in Seq2Seq models. The last layer is a feed-forward neural network, similar to the Encoder layer.
@@
The most common use of the Transformer is as the backbone for a fine-tuned pre-trained transformer-based language model, such as BERT [15], described in subsection 4.10. These language models are trained with large amounts of data to perform many tasks and only need to be fine-tuned for the specific task at hand. In the case of a classification task such as Emotion Recognition, one just needs to add a linear layer or a more complex neural network architecture on top of the language model. The training is performed with supervised learning, resorting to backpropagation algorithms.
@@ 4.9.1 Knowledge-Enriched Transformer [144]
The invention of the Transformer [112] led to new state-of-the-art results in several Natural Language Processing tasks. The Knowledge-Enriched Transformer (KET) uses self-attention to model context and response using the Transformer, that a has shorter path of information flow than gated RNNs, overcoming the difficulty in capturing long-term dependencies. Conversations are modeled in the entire Transformer as a single input.
First, concepts are retrieved for each word in the conversation using an external knowledge base, a graph of concepts, then the concept representation is computed using a dynamic context-aware graph attention mechanism over the concepts that are then combined with the input sentence embeddings. Self-attention is used to learn utterance representations individually and hierarchical self-attention is applied to the context to learn the context representation. Finally, the encoder and decoder attention mechanism is applied, followed by max pooling and a linear layer. KET achieves an F1-score of 59.56% on the IEMOCAP dataset.
@@ 4.9.3 DialogXL [98]
DialogXL is an adaptation of the pre-trained language model XLNet [130], which is based on the Transformer-XL [11], for Emotion Recognition in Conversations. It replaces XLNet’s segment recurrence with a memory utterance recurrence to leverage historical utterances. These utterances’ hidden states are stored in a memory bank to reuse them while identifying a query utterance. The approach also replaces XLNet’s vanilla self-attention proposing a dialog-aware self-attention to grasp useful intra- and inter-speaker dependencies. This is composed of four types of self-attention: global and local self-attention for different sizes of receptive fields, and speaker and listener self-attention for intra- and inter-speaker dependencies. DialogXL achieves an F1-score of 65.94% on the IEMOCAP dataset.
4.10 Embeddings from Pre-Trained Transformer Language Models
@@
Static embeddings [74, 80, 75] are obtained based on the co-occurrence of adjacent words and have a fixed representation for the words, not taking into account their context.
Contrarily, contextual embeddings include information from the context in the word representation, a representation that differs according to the occurrence of the word. The only disadvantage of this kind of representation is that we cannot use pre-trained embeddings as in static embeddings so there is an inference time for generating the embeddings. The models are pre-trained with large amounts of data, in order to enable generating good embedings. A first approach, Embeddings from Language Models (ELMo) [83] is based on a stack of two Bi-LSTMs, leveraging the full context of a word. It provides a context-free representation of the word along with context information of the sense of the word and its syntax. The second approach, Bidirectional Encoder Representations from Transformers (BERT) [15], built from a Transformer encoder stack, leads to state-of-the-art results in multiple Natural Language Processing tasks. The input given to BERT is the sum of three embeddings, the token embedding corresponding to the sequence of words, the segment embedding containing information on which sentence each token belongs to, and the position embedding, defining the position of each token in the sequence and distance to other tokens taking into account the order of the sequence. An improvement upon BERT, RoBERTa [65] is trained with more data and for a longer period of time. Given that context is key for text Emotion Recognition in Conversations, contextual word embeddings are more adequate for the task than static embeddings, as can be seen in the higher performance of works that resort to these embeddings, namely resultant from Pre-trained language models, such as BERT-based embeddings. The latter are also more tailored to deal with common sense, informal language, and sarcasm, ERC challenges described in Section 2, since they are pre-trained on large amounts of text.
@@ 4.10.1 COSMIC [25]
COSMIC extracts commonsense features resorting to COMET [4], a commonsense transformer encoder-decoder model trained on the task of generative commonsense knowledge construction that uses GPT [93] as its generative model. The approach uses RoBERTa for context-independent feature extraction, passes each utterance through COMET’s encoder, and extracts the activations from the final time step. COSMIC maintains a context state and attention vector which are always shared between the participants of the conversation. Five GRUs are used to model context state, internal state, external state, intent state, and emotion state, being the latter used for emotion classification. This work achieves an F1-score of 65.28% in the IEMOCAP dataset.
@@ 4.10.2 Psychological [58]
Psychological also leverages COMET as the commonsense knowledge base and resorts to RoBERTa as the utterance-level encoder. It proposes SKAIG as a conversation-level encoder, a locally connected graph, where the targeted utterance receives information from that past and future context and is also self-connected.
Assuming that the influence of an utterance on contextual utterances is locally effective, the targeted node is connected with contextual nodes in a window of a given size of utterances in the past and future. Knowledge from COMET is introduced to enrich the edges with different relations. To propagate information through SKAIG a Graph Transformer [101] is used. Finally, a linear unit predicts the emotion distributions. This work achieves an F1-score of 66.96% in the IEMOCAP dataset.
@@ 4.10.3 DAG-ERC [99]
Graph-based methods tend to neglect distant utterances and sequential information. On the other hand, recurrence-based methods leverage those but tend to update the query utterance’s state with limited information from the nearest utterances. Combining both graph and recurrence structures explores their advantages and mitigates their drawbacks. This is the idea behind DAG-ERC, which regards each conversation as a directed acyclic graph (DAG), a combination of both structures. DAG-ERC is based on the DAGNN [105] architecture, with two improvements: a relation-aware feature transformation that gathers information based on speaker identity and a contextual information unit that enhances the information of historical context. RoBERTa-Large is used as the feature extractor. It achieves an F1-score of 68.03% on the IEMOCAP dataset, outperforming the previous approaches.
4.11 ERC within a Generative Framework
With the advent of Generative Large Language Models (LLMs), such as the GPT series [92, 93, 5] and Llama [106], classification tasks started to be reformulated within generative frameworks. The idea is to prompt the LLM with the utterance to be classified and other relevant additional input and request an emotion label for the utterance. Open-source LLMs can also be fine-tuned and even pre-trained.
@@ 4.11.1 InstructERC [54]
In InstructERC the prompt consists of the utterance to be classified, the ERC instruction, the conversational context, the set of possible classification labels, and similar utterances together with its emotion labels. The LLM is pre-trained with a speaker identification task and fine-tuned both with the main ERC task and an emotion influence prediction task to improve overall performance. Experiments were made with several LLMs, and the best results were obtained with LLama2. InstructERC achieves 71.39% in the IEMOCAP dataset, outperforming the previous approaches.
@@ 5 Advisable ERC Practices
This section provides useful methods for better ERC frameworks, from different ways of dealing with subjectivity in emotion annotation and modelling, to several ways of dealing with the typically unbalanced ERC datasets.
@@ 5.1 Learning Subjectivity
All human annotation tasks involve some degree of uncertainty, that can stem amongst other factors from the subjectivity inherent to the annotator. Since Emotion Recognition is amongst the most subjective tasks [111] and that subjectivity is the main source of uncertainty in Emotion Recognition annotation, we address uncertainty as subjectivity in this survey [95]. A good annotation practice is to resort to several annotators to avoid biasing the results towards a single annotator’s subjectivity. This will naturally yield different labels from each annotator, the so called annotation disagreement. The more appropriately one deals with the disagreement in the annotated labels the more reliable will the data and classifiers be. Several techniques for dealing with disagreement are described in this section. They are divided into techniques that address subjectivity by adapting the labels, towards an estimate of a gold label, and techniques that embrace subjectivity by adapting the classifier architectures to deal with the subjectivity of the labels, leveraging such property.
@@ 5.1.1 Addressing Subjectivity
For categorical labels, considering that a gold or true label exists for each sample, the simple way to obtain it would be through majority voting, and choosing the label with the most annotations. However, this does not account for the different annotator skill levels or sample inherent difficulty.
@@
For interval labels, weighted fusion of data can be performed with approaches such as the Evaluator Weighted Estimator (EWE) [28], which weights each rater’s annotations based on its rater-specific inter-reliability scores, or the Weighted Trustability Evaluator (WTE) [29], which is instead based on each rater’s performance consistency.
@@ 5.1.2 Embracing Subjectivity [95]
The annotator (dis)agreement can be used as privileged information for master-class learning, being viewed as samples’ additional information [21] to facilitate learning in the training phase, that might or might not be required for making predictions during testing, for example, to weigh positively the loss of the corresponding samples. In this setting, it is useful to learn the meaning of high-rater disagreement for the particular dataset in use.
Instead of considering the existence of a hard label, as in subsubsection 5.1.1, one can calculate the distribution of label annotations to define a soft label distribution per sample, which encodes the subjectivity for each sample and allows for the classifier to learn label correlations. Another approach is to model the annotations of each particular rater, through an ensemble of models. It allows to give more importance to expert raters rather than novices or spammers. The addition of an "unsure" label would facilitate the annotator’s workload and also make the model learn the ambiguity ground-truth, although not having to predict "unsure" labels during testing. Furthermore, methods to understand whether a sample is mislabelled by certain raters or inherently ambiguous could inform the model and annotators during active learning processes.